# access point ADAS advanced driver-assistance systems

### v1

@@
Target Tracking: Statistics of Successive Successful Target Detection in Automotive Radar Networks
Gourab Ghatak, Member, IEEE The author is with the Department of Electrical Engineering at the Indian Institute of Technology (IIT) Delhi, New Delhi, India 110016. Email: gghatak@ee.iitd.ac.in.
@@ Abstract
Index Terms:
Radar tracking, automotive radars, stochastic geometry, vehicular networks.
@@ I Introduction
Automotive radars play a critical role in modern advanced driver-assistance systems (ADAS) and autonomous driving technologies. These systems rely on accurate detection and tracking of moving objects such as vehicles, pedestrians, and other obstacles [1, 2]. In particular, radar-based target tracking ensures timely and accurate information about the target’s position, velocity, and trajectory, allowing for safer navigation and collision avoidance [3]. Given the inherent stochastic nature of the wireless environment, especially in densely populated urban scenarios, developing robust radar tracking algorithms is essential. Radar tracking is a two-step process, comprising both detection and subsequent tracking over multiple time slots with consecutive radar pulses [4]. The first step involves the successful detection of the target based on the reflected signal signal-to-interference-plus-noise ratio (SINR[5]. In a given time slot, the target is successfully detected if the SINR exceeds a predefined threshold. Then, a target is considered successfully tracked if it is detected successfully in $\nu$ consecutive slots.
@@ I-A Motivation
The spatial statistics of the detection performance has been thoroughly investigated using tools from stochastic geometry, e.g., see [6, 5]. In particular, considering the vehicular node locations as points of a point process, researchers have derived the probability of successful detection, the meta-distribution of the reflected signal SINR, and consequently, the mean local delay in detection. However, the statistics of consecutive successful detection events is unexplored, which we address in this work. Additionally, our model accounts for street intersections and the corresponding incoming traffic in analyzing the system performance. This is critical for developing automotive radar aided safety applications but has limited literature.
@@ I-B Related Work
In recent years, studies based on stochastic geometry have gained traction to model and analyze the statistics of automotive radar interference while addressing the variability in radar deployments, e.g., see [6, 7, 8, 5, 9]. These models are crucial for accurately evaluating radar system performance and determining the impact of interference mitigation strategies. For example, the study in [6] modeled the distribution of automotive radars using a Poisson point process (PPP) and estimated the average signal-to-interference ratio (SIR) for radar detection. Munari et al. [7] employed the strongest interferer approximation to assess radar detection range and false alarm rates. Additionally, the work in [8] examined radar detection probabilities by incorporating fluctuating radar cross-sections (RCS) based on Swerling-I and Chi-square models, offering more precise and intuitive outcomes compared to constant RCS models. A fine grained analysis of the impact of interference is studied in [5] where the authors derived not only the meta distribution (MD) of the SINR but also employed it to study the mean local delay in radar detection. Furthermore, they proposed an online algorithm to control the channel access by individual radars. Radar medium access control (MAC) protocols such as ALOHA and carrier sense multiple access (CSMA) have also been investigated in [10, 11]. Furthermore, asynchronous non-cooperative protocols for MAC was studied in [12]. By enhancing these models, we can improve performance evaluations, leading to the development of more optimized solutions. One of the primary advantages of using stochastic geometry frameworks is the ability to explore various spatial scenarios with significantly reduced computational resources and time compared to traditional system-level Monte Carlo simulations or real-world measurement data collection.
@@
While these studies predominantly focus on two-lane scenarios, the authors in [13] extended the analysis to multi-lane scenarios by utilizing a marked point process model to characterize radar interference. Furthermore, more recent work has incorporated Matern Hard-Core Processes (MHCP) [14, 15], which account for vehicle dimensions and the required road space, providing a more realistic interference model for automotive radar systems. However, in order to gain a more accurate view of the network, a complete two dimensional spatial model is paramount, which none of the above works address. In this regard, novel line processes have been proposed to model and study structure of streets and vehicular networks [16, 17]. In this work, we consider the Poisson line Cox process (PLCP) model for emulating the streets in order to accurately incorporate the impact of radar beamwidth (e.g., see [18] for factors impacting beamwidth selection in such networks) and the incoming traffic at intersections.
@@
More importantly, all the above studies either focus on one shot detection, i.e., the event of successful target detection in one radar pulse transmission attempt, or characterize the mean local delay, i.e., the number of transmission attempts for first successful detection. Albeit useful, these metrics cannot estimate the performance of target tracking. To address this, we derive the distribution of run-length of successes and accordingly, characterize the statistics of successive successful detection, or tracking [^1] From a control systems perspective, the run-length distribution was also characterized in [19], however, there the authors did not study the same in a PLCP model and did not employ the same to study an automotive radar network, which presents its own challenges..
@@ I-C Contributions
  • Then, we introduce a novel block ALOHA-based radar MAC protocol where active radars access the channel in a block of $T$ consecutive slots with block probability $\delta$. Unlike the slot-based ALOHA protocol typically assumed in automotive radar networks (e.g., see [5, 7]), the block ALOHA enables an enhanced radar tracking probability. Then we derive the detection probability of the ego radar, followed by the conditional distribution of the number of successive slots in which the reflected signal SINR exceeds a detection threshold. Furthermore, we derive the expected number of such successful tracking events in the block. As deriving the distribution of the conditional tracking probability is infeasible, we analyze its first moment, which denotes the probability of successful target tracking across all network realizations.
@@
The rest of the paper is organized as follows. In Section II we describe the system model, characterize the interference set, and outline the key definitions of radar detection and radar tracking. Our main results are presented in Section III where we derive the statistics of radar tracking. Then, in Section IV, based on the derived framework, we explore QoS-Aware optimal frame design. Numerical results to highlight the salient features and insights in our framework are presented in Section V. Finally, the paper concludes in Section VI.
@@ II-A Street Geometry and Vehicle Locations
We consider a network of streets modeled as a homogeneous Poisson line process (PLP), $\mathcal{P}=\{L_{1},L_{2},\dots\}$ [20]. A line $L_{i}\in\mathcal{P}$ corresponds to a tuple $(\theta_{i},r_{i})$, where $r_{i}$ is the length of the normal to $L_{i}$ from the origin and $\theta_{i}$ is the angle between the normal and the $x$-axis. The parameters $\mathbf{q}_{i}\coloneq(\theta_{i},r_{i})$ reside in the representation space $\mathcal{D}\equiv[0,2\pi)\times(0,\infty)$. The number of parameter points, $I$, in any $S\subset\mathcal{D}$ follows the Poisson distribution with parameter $\lambda_{\rm L}|S|$, where $|S|$ represents the Lebesgue measure of $S$.
The locations of the vehicles with mounted radars on the street $L_{i}$ follow a one-dimensional PPP, $\Phi_{{\rm L}_{i}}$, with intensity $\lambda$, that represents the vehicular density. The PPP of the vehicle locations on any street is independent of the corresponding distributions on the other streets. Hence, the complete distribution of the vehicles is a homogeneous PLCP $\Phi$, on the domain $\mathcal{P}$ where $\Phi=\bigcup_{{\rm L}_{i}\in\mathcal{P}}\Phi_{{\rm L}_{i}}$. Our analysis is carried out from the perspective of an ego radar located at the origin of the two-dimensional plane. We model the ego radar as the typical point of $\Phi$. In particular, let us define
@@
for any point $o$, and note that under expectation over $\Phi^{o}$, $o$ becomes the typical point of the process $\Phi$ [21]. Since by construction, $\Phi$ is a stationary process, let us consider $o$ to be the origin of $\mathbb{R}^{2}$. Under Palm conditioning, a line $L_{0}$ passes through the origin almost surely, which without loss of generality, we consider to be the $y-$axis. The ego radar experiences interference if and only if both the ego radar and interfering radar fall into each other’s radar sectors simultaneously. To model the interference from the vehicles traveling to the opposite side as the ego radar, we consider on line $L_{0}$ the location of vehicles follows a one-dimensional PPP having the same intensity as the remaining lines. Therefore, the overall spatial stochastic process of potential interfering radars, $\Phi_{{\rm P}_{0}}$, as per the Palm conditioning, is the superposition of $\Phi$ and an independent 1D PPP on line $L_{0}$, i.e., $\Phi_{{\rm P}_{0}}=\Phi\cup\Phi_{L_{0}}$.
@@ II-B Signal and Channel Model
The radars have a half-power beamwidth of $\Omega$, and the target is assumed to be located at a distance $R$ from the radar on the same street within the main beam of the ego radar. Let the transmit power be $P$ and the path-loss exponent be $\alpha$. The target is assumed to have a Swerling-I fluctuating radar cross-section, $\sigma_{\rm c}$, with a mean of $\bar{\sigma}$ [22]. Let $G_{t}$ be the gain of the transmitting antenna and $A_{\rm e}$ the effective area of the receiving antenna aperture. Then, the reflected power from the target at the ego radar is $S=\gamma\sigma_{\rm c}PR^{-2\alpha}$, where $\gamma=\frac{G_{\rm t}}{(4\pi)^{2}}A_{\rm e}$. Let the coordinates of any interfering radar be denoted by $\mathbf{w}=(x,y)$, such that $x\cos\theta_{i}+y\sin\theta_{i}=r_{i}$ for $(\theta_{i},r_{i})\in\mathcal{D}_{\rm P}$. The interfering signals undergo independent multi-path Rayleigh fading with parameter 1 [5]. Thus, the interference at the ego radar due to an interfering radar located at $w$ is $\mathbf{I}=P\gamma h_{\mathbf{w}}||\mathbf{w}||^{-\alpha}$, where $h_{\mathbf{w}}$ is the fading power. Now, assuming that all the automotive radars share the same power and gain characteristics, the SINR at the ego radar is
$\xi(R)=\frac{\gamma\sigma_{\rm c}PR^{-2\alpha}}{N_{0}+\sum_{\mathbf{w}\in\Phi_{\rm I}}4\pi\gamma Ph_{\mathbf{w}}||\mathbf{w}||^{-\alpha}},$ (1)
where $\Phi_{\rm I}\subset\Phi_{{\rm P}_{0}}$ is the interfering set, defined and derived in the next subsection.
@@ II-C Interfering Set and Distance
The orientation of the vehicles on the street depends on the generating angle of the street and the direction of their movement. Thus, the boresight direction of any radar on the line $L_{i}$ is given by the two unit vectors: $\mathbf{z}_{i}$ and $-\mathbf{z}_{i}$, where $\mathbf{z}_{i}=(-\sin\theta_{i},\cos\theta_{i})^{T}$. The parameters $(\theta_{i},r_{i})$, determines the direction of the radar sector on the street $L_{i}$. Any point $(p,q)\in\mathbb{R}^{2}$ lies in the sector if the angle made by the displacement vector between $(p,q)$ and $(x,y)$ and the boresight direction is larger than $\cos\Omega$. Thus, the sector of a radar located at $(x,y)$ is either $\mathcal{R}^{+}_{(x,y)}$ or $\mathcal{R}^{-}_{(x,y),k}$ based on its direction, mathematically defined below.
@@
[17] The radar sector of a vehicle located at $(x,y)\in\Phi_{{\rm P}_{0}}$ for boresight direction $\mathbf{z}_{i}$ is
$\mathcal{R}^{+}_{(x,y)}$ =$=\Bigg{\{}(p,q)\in\mathbb{R}^{2}\colon\frac{\big{(}(p,q)-(x,y)\big{)}\cdot\mathbf{z}_{i}}{||(p,q)-(x,y)||}>\cos\Omega,$
@@
=$=\Bigg{\{}(p,q)\in\mathbb{R}^{2}\colon\frac{\cos\theta_{i}\left(q-y\right)-\sin\theta_{i}\left(p-x\right)}{\sqrt{(p-x)^{2}+(q-y)^{2}}}>\cos\Omega,$
@@
And for boresight direction $-\mathbf{z}_{i}$,
@@
$\Bigg{\{}(p,q)\in\mathbb{R}^{2}\colon\frac{\sin\theta_{i}\left(p-x\right)-\cos\theta_{i}\left(q-y\right)}{\sqrt{(p-x)^{2}+(q-y)^{2}}}>\cos\Omega,$
@@
The ego radar sector corresponds to the line $(\theta_{i},r_{i})=(0,0)$ and the location $(x,y)=(0,0)$. If ego radar is located within the radar sector of the radar present at $(x,y)$ i.e., $\mathbf{1}\left((0,r)\in\mathcal{R}^{+}_{(x,y)}\right)+\mathbf{1}\left((0,r)\in\mathcal{R}^{-}_{(x,y)}\right)$ evaluates to one, and simultaneously the automotive radar located at $(x,y)$ is located in the interior region of ego radar sector i.e., $\mathbf{1}\left((x,y)\in\mathcal{R}^{+}_{(0,r)}\right)+\mathbf{1}\left((x,y)\in\mathcal{R}^{-}_{(0,r)}\right)$ evaluates to one, then radar at $(x,y)$ will be an element of $\Phi_{\rm I}$. From the perspective of the ego radar, the following result characterizes the location of the interfering radars on each street $L_{i}$.
@@
$b$ $=\frac{\sin\Omega/2}{d\sin\left(\theta-\Omega/2\right)}.$

### v2

@@
URLLC
ultra-reliable low-latency communications
Statistics of Successive Successful Target Detection in Automotive Radar Networks
Gourab Ghatak, Member, IEEE The author is with the Department of Electrical Engineering at the Indian Institute of Technology (IIT) Delhi, New Delhi, India 110016. Email: gghatak@$e$.iitd.ac.in.
@@ I Introduction
Automotive radars play a critical role in modern advanced driver-assistance systems (ADAS) and autonomous driving technologies. These systems rely on accurate detection and tracking of moving objects such as vehicles, pedestrians, and other obstacles [dong2020probabilistic, patole2017automotive]. In particular, radar-based target tracking ensures timely and accurate information about the target’s position, velocity, and trajectory, allowing for safer navigation and collision avoidance [cao2021automotive]. Given the inherent stochastic nature of the wireless environment, especially in densely populated urban scenarios, developing robust radar tracking algorithms is essential. Radar tracking is a two-step process, comprising both detection and subsequent tracking over multiple time slots with consecutive radar pulses [howard1990tracking]. The first step involves the successful detection of the target based on the reflected signal signal-to-interference-plus-noise ratio (SINR[ghatak2022radar]. In a given time slot, the target is successfully detected if the SINR exceeds a predefined threshold. Then, a target is considered successfully tracked if it is detected successfully in $\nu$ consecutive slots.
@@ I-A Motivation
The spatial statistics of the detection performance has been thoroughly investigated using tools from stochastic geometry, e.g., see [al2017stochastic, ghatak2022radar]. In particular, considering the vehicular node locations as points of a point process, researchers have derived the probability of successful detection, the meta-distribution of the reflected signal SINR, and consequently, the mean local delay in detection. However, the statistics of consecutive successful detection events is unexplored, which we address in this work. Additionally, our model accounts for street intersections and the corresponding incoming traffic in analyzing the system performance. This is critical for developing automotive radar aided safety applications but has limited literature.
@@ I-B Related Work
In recent years, studies based on stochastic geometry have gained traction to model and analyze the statistics of automotive radar interference while addressing the variability in radar deployments, e.g., see [al2017stochastic, munari2018stochastic, fang2020stochastic, ghatak2022radar, ghatak2021fine]. These models are crucial for accurately evaluating radar system performance and determining the impact of interference mitigation strategies. For example, the study in [al2017stochastic] modeled the distribution of automotive radars using a Poisson point process (PPP) and estimated the average signal-to-interference ratio (SIR) for radar detection. Munari et al. [munari2018stochastic] employed the strongest interferer approximation to assess radar detection range and false alarm rates. Additionally, the work in [fang2020stochastic] examined radar detection probabilities by incorporating fluctuating radar cross-sections (RCS) based on Swerling-I and Chi-square models, offering more precise and intuitive outcomes compared to constant RCS models. A fine grained analysis of the impact of interference is studied in [ghatak2022radar] where the authors derived not only the meta distribution (MD) of the SINR but also employed it to study the mean local delay in radar detection. Furthermore, they proposed an online algorithm to control the channel access by individual radars. Radar medium access control (MAC) protocols such as ALOHA and carrier sense multiple access (CSMA) have also been investigated in [haritha2021slotted, sukumaran2022slotted]. Furthermore, asynchronous non-cooperative protocols for MAC was studied in [9399786]. By enhancing these models, we can improve performance evaluations, leading to the development of more optimized solutions. One of the primary advantages of using stochastic geometry frameworks is the ability to explore various spatial scenarios with significantly reduced computational resources and time compared to traditional system-level Monte Carlo simulations or real-world measurement data collection. Additionally, it must be noted that while real-world radar datasets and high fidelity system-level simulations are valuable for empirical benchmarking, they remain difficult to scale or generalize due to their inherent limitations in capturing large-scale interference and deployment randomness. In contrast, our stochastic geometry framework balances realism and analytical tractability, enabling scalable and generalizable performance analysis.
@@
While these studies predominantly focus on two-lane scenarios, the authors in [chu2020interference] extended the analysis to multi-lane scenarios by utilizing a marked point process model to characterize radar interference. Furthermore, more recent work has incorporated Matern Hard-Core Processes (MHCP) [mishra2020stochastic, kui2021interference], which account for vehicle dimensions and the required road space, providing a more realistic interference model for automotive radar systems. However, in order to gain a more accurate view of the network, a complete two dimensional spatial model is paramount, which none of the above works address. In this regard, novel line processes have been proposed to model and study structure of streets and vehicular networks [shah2024binomial, shah2024modeling]. In this work, we consider the Poisson line Cox process (PLCP) model for emulating the streets in order to accurately incorporate the impact of radar beamwidth (e.g., see [nabil2024beamwidth] for factors impacting beamwidth selection in such networks) and the incoming traffic at intersections.
@@
More importantly, all the above studies either focus on one shot detection, i.e., the event of successful target detection in one radar pulse transmission attempt, or characterize the mean local delay, i.e., the number of transmission attempts for first successful detection. Albeit useful, these metrics cannot estimate the performance of target tracking. To address this, we derive the distribution of run-length of successes and accordingly, characterize the statistics of successive successful detection, or tracking [^1] From a control systems perspective, the run-length distribution was also characterized in [ghatak2024RAWNET], however, there the authors did not study the same in a PLCP model and did not employ the same to study an automotive radar network, which presents its own challenges..
@@ I-C Contributions
  • Then, we introduce a novel block ALOHA-based radar MAC protocol where active radars access the channel in a block of $T$ consecutive slots with block probability $\delta$. Unlike the slot-based ALOHA protocol typically assumed in automotive radar networks (e.g., see [ghatak2022radar, munari2018stochastic]), the block ALOHA enables an enhanced radar tracking probability. Then we derive the detection probability of the ego radar, followed by the conditional distribution of the number of successive slots in which the reflected signal SINR exceeds a detection threshold. Furthermore, we derive the expected number of such successful tracking events in the block. As deriving the distribution of the conditional tracking probability is infeasible, we analyze its first moment, which denotes the probability of successful target tracking across all network realizations.
@@
I-D Key Technical Challenges:
Most of the stochastic-geometry based analysis of automotive radar networks were focused on a single or double lane highway scenario with a simple one dimensional PPP to emulate the location of the interferers (e.g., see [al2017stochastic]). Although the PLCP is a well-studied component in stochastic-geometry, most of its characterization has focused on the perspective of cellular networks. On the contrary, due to the consideration of automotive radar networks in this paper, the interfering set of the vehicles is formed by the intersection of the PLCP with a double conic set. This makes the interfering set a non-stationary point process, which requires the derivation of the conditional interference lengths for a given intersection distance. The challenge is compounded by the fact that in contrast to the typical one-shot detection probability, due to the correlation of run-lengths, the tracking characterization is a challenging task.
The rest of the paper is organized as follows. In Section I we describe the system model, characterize the interference set, and outline the key definitions of radar detection and radar tracking. Our main results are presented in Section III where we derive the statistics of radar tracking. Then, in Section IV, based on the derived framework, we explore QoS-Aware optimal frame design. Numerical results to highlight the salient features and insights in our framework are presented in Section V. Finally, the paper concludes in Section VI.
@@ II System Model and Key Definitions
Parameter name Symbol Units
Transmit power P dBW
Target distance R m
Radar cross section $\sigma_{\rm c}$ m2
Path loss exponent $\alpha$ -
Noise power $N_{0}$ dBW
SINR Threshold $\gamma$ dB
SINR $\xi$ dB
Tracking length $L$ slots
Reliability (minimum tracking run length) $\nu$ slots
Block length T slots
Minimum and maximum interference region in an intersecting street $a$ and $b$ m
TABLE I: System parameters and definitions.
@@ II-A Street Geometry and Vehicle Locations
We consider a network of streets modeled as a homogeneous Poisson line process (PLP), $\mathcal{P}=\{L_{1},L_{2},\dots\}$ [ghatak2019small]. A line $L_{i}\in\mathcal{P}$ corresponds to a tuple $(\theta_{i},r_{i})$, where $r_{i}$ is the length of the normal to $L_{i}$ from the origin and $\theta_{i}$ is the angle between the normal and the $x$-axis. The parameters $\mathbf{q}_{i}:=(\theta_{i},r_{i})$ reside in the representation space $\mathcal{D}\equiv[0,2\pi)\times(0,\infty)$. The number of parameter points, $I$, in any $S\subset\mathcal{D}$ follows the Poisson distribution with parameter $\lambda_{\rm L}|S|$, where $|S|$ represents the Lebesgue measure of $S$. The locations of the vehicles with mounted radars on the street $L_{i}$ follow a one-dimensional PPP, $\Phi_{{\rm L}_{i}}$, with intensity $\lambda$, that represents the vehicular density. The PPP of the vehicle locations on any street is independent of the corresponding distributions on the other streets. Hence, the complete distribution of the vehicles is a homogeneous PLCP $\Phi$, on the domain $\mathcal{P}$ where $\Phi=\bigcup_{{\rm L}_{i}\in\mathcal{P}}\Phi_{{\rm L}_{i}}$. Our analysis is carried out from the perspective of an ego radar located at the origin of the two-dimensional plane. We model the ego radar as the typical point of $\Phi$. It is important of highlight that while this assumption facilitates tractable analysis, it does not capture potential correlations in the spatial deployment of radars along structured line processes (e.g., roadways or corridors). An alternative approach is to employ the binomial line Cox process (BLCP) or inhomogeneous PLCP, which allows for more realistic modeling of radar interferers distributed along random lines. Incorporating such inhomogeneity is left outside the present scope and is discussed in detail in the numerical results section. In particular, let us define
@@
for any point $o$, and note that under expectation over $\Phi^{o}$, $o$ becomes the typical point of the process $\Phi$ [chiu2013stochastic]. Since by construction, $\Phi$ is a stationary process, let us consider $o$ to be the origin of $\mathbb{R}^{2}$. Under Palm conditioning, a line $L_{0}$ passes through the origin almost surely, which without loss of generality, we consider to be the $y$-axis. The ego radar experiences interference if and only if both the ego radar and interfering radar fall into each other’s radar sectors simultaneously. To model the interference from the vehicles traveling to the opposite side as the ego radar, we consider on line $L_{0}$ the location of vehicles follows a one-dimensional PPP having the same intensity as the remaining lines. Therefore, the overall spatial stochastic process of potential interfering radars, $\Phi_{{\rm P}_{0}}$, as per the Palm conditioning, is the superposition of $\Phi$ and an independent 1D PPP on line $L_{0}$, i.e., $\Phi_{{\rm P}_{0}}=\Phi\cup\Phi_{L_{0}}$.
@@ II-B Signal and Channel Model
[…]
where where $N_{0}$ is the noise power, i.e., the noise density multiplied by the transmission bandwidth and $\Phi_{\rm I}\subset\Phi_{{\rm P}_{0}}$ is the interfering set, defined and derived in the next subsection.
@@ II-C Interfering Set and Distance
The orientation of the vehicles on the street depends on the generating angle of the street and the direction of their movement. Thus, the boresight direction of any radar on the line $L_{i}$ is given by the two unit vectors: $\mathbf{z}_{i}$ and $\mathbf{z}_{i}$, where $\mathbf{z}_{i}=(-\sin\theta_{i},\cos\theta_{i})^{T}$. The parameters $(\theta_{i},r_{i})$, determines the direction of the radar sector on the street $L_{i}$. Any point $(p,q)\in\mathbb{R}^{2}$ lies in the sector if the angle made by the displacement vector between $(p,q)$ and $(x,y)$ and the boresight direction is larger than $\cos\Omega$. Thus, the sector of a radar located at $(x,y)$ is either $\mathcal{R}^{+}_{(x,y)}$ or $\mathcal{R}^{-}_{(x,y),k}$ based on its direction, mathematically defined below.
@@
[shah2024modeling] The radar sector of a vehicle located at $(x,y)\in\Phi_{{\rm P}_{0}}$ for boresight direction $\mathbf{z}_{i}$ is
[…]
@@
$\hskip-35.56593pt=\Bigg\{(p,q)\in\mathbb{R}^{2}\colon\frac{\cos\theta_{i}\left(q-y\right)-\sin\theta_{i}\left(p-x\right)}{\sqrt{(p-x)^{2}+(q-y)^{2}}}>\cos\Omega,$
@@
And for boresight direction $\mathbf{z}_{i}$,
@@
$\hskip-28.45274pt\Bigg\{(p,q)\in\mathbb{R}^{2}\colon\frac{\sin\theta_{i}\left(p-x\right)-\cos\theta_{i}\left(q-y\right)}{\sqrt{(p-x)^{2}+(q-y)^{2}}}>\cos\Omega,$
@@
Figure 1: Illustration of the interfering ranges in an intersecting street in $L_{0}$.
The ego radar sector corresponds to the line $(\theta_{i},r_{i})=(0,0)$ and the location $(x,y)=(0,0)$. If ego radar is located within the radar sector of the radar present at $(x,y)$ i.e., $\mathbf{1}\left((0,r)\in\mathcal{R}^{+}_{(x,y)}\right)+\mathbf{1}\left((0,r)\in\mathcal{R}^{-}_{(x,y)}\right)$ evaluates to one, and simultaneously the automotive radar located at $(x,y)$ is located in the interior region of ego radar sector i.e., $\mathbf{1}\left((x,y)\in\mathcal{R}^{+}_{(0,r)}\right)+\mathbf{1}\left((x,y)\in\mathcal{R}^{-}_{(0,r)}\right)$ evaluates to one, then radar at $(x,y)$ will be an element of $\Phi_{\rm I}$. Mathematically, this thinning procedure is represented as
$\Phi_{\rm I}=\Phi_{{\rm P}_{0}}\cap\mathcal{DC}(r),$ (2)
where $\mathcal{DC}(r)$ is the double-conic intersection region represented as:
$\mathcal{DC}(r)=$ $\{(x,y):\left[\mathbf{1}\left((0,r)\in\mathcal{R}^{+}_{(x,y)}\right)+\mathbf{1}\left((0,r)\in\mathcal{R}^{-}_{(x,y)}\right)\right]$
$\cdot\left[\mathbf{1}\left((x,y)\in\mathcal{R}^{+}_{(0,r)}\right)+\mathbf{1}\left((x,y)\in\mathcal{R}^{-}_{(0,r)}\right)\right]=1\}.$ (3)
For the ego-radar, $\Phi_{\rm I}$ represents the set of vehicles which lie in the ego-radar’s beam while simultaneously the ego-radar lies in their beam, thereby causing mutual interference. This is further illustrated in Fig. 1, where the ego radar is shown in green. An intersecting street $L_{i}$ crosses $L_{0}$ at a distance $d_{i}$ from the ego radar. As a result, the variables $a$ and $b$ are defined from the point of intersection of $L_{i}$ and $L_{0}$. In order to calculate $w$, we further need the angle of intersection, which is denoted by $\theta_{i}$ in the figure. Two extremes of interfering radar locations are shown with blue (at location $T$) and magenta (in location $S$), respectively. Now, on this street $L_{i}$, any vehicle if present between $S$ and $T$ will create an interference to the ego-radar. Any vehicle which is present at a distance greater than $b$ (i.e., behind $S$) will not create interference since it will not be present in the ego-radar beam. On the contrary, any radar present at a distance lower than $a$ (i.e., in front of $T$) also will not cause interference to the ego radar since the ego-radar will not be present in the beam of such a vehicle. For all such streets intersecting $L_{0}$, the interfering region needs to be characterized for such $a$ and $b$ limits. This is formally presented in the following result.
@@
$b$ $=\frac{d\sin\Omega/2}{\sin\left(\theta-\Omega/2\right)}.$